| Name | Email | Date | Comment
|
| ANS
| anolla@naic.edu
| 8/4/2000
| These receiver was installed during July , 2000.It was calibrated on 8/3/2000 ,the
cal values are Chn. A 186 k deg & Chn B is a 150 K deg.
|
| ANS
| anolla@naic.edu
| 8/25/2000
| Dewar taken down for repairs, it is not holding the vacuum.
|
| ANS
| anolla@naic.edu
| 9/20/2000
| Dewar was taken up on 9/18/2000.It was tested in the cryolab for more than a
week. Coco found a couple of leaks at the Dewar; also he replaced the crosshead on it.
New Cal & RF cables were done
|
| ANS
| anolla@naic.edu
| 10/10/2000
| Receiver warmed up due to power dip on Saturday that tripped off the compressor.
Was cooled down.
|
| ANS
| anolla@naic.edu
| 10/11/2000
| There is a problem with the receiver cryo and it looks like if the compressor is not
holding the 3 crossheads attached to it.
|
| ANS
| anolla@naic.edu
| 11/8/2000
| Start cooling down using the Austin small cryo comp.
|
| ANS
| anolla@naic.edu
| 11/9/2000
| Dewar cooled , cryo temp is 49.2 ºk. Austin comp 125 used.
|
| ANS
| anolla@naic.edu
| 11/14/2000
| Dewar at room temp due that was necessary to use the Austin 125 (small) cryo
comp. to cool down the 430 Mhz Dewar.
|
| ANS
| anolla@naic.edu
| 12/13/2000
| Channel B RF out was not connected .ANS hooked up a 327 Mhz filter on it.
|
| ANS
| anolla@naic.edu
| 12/15/2000
| JR removed the matching xformers & hybrids out of the antenna.
|
| ANS
| anolla@naic.edu
| 12/28/2000
| JR move the receiver to the new open frame cabinet at the rotary floor.
|
| ANS
| anolla@naic.edu
| 1/31/2001
| Receiver cooled with comp. 008.
|
| ANS
| anolla@naic.edu
| 2/8/2001
| The A/L power supply was not working is suspected that the DS chip is burned
out.
|
| ANS
| anolla@naic.edu
| 3/12/2001
| Dewar hot due to cryo comp shutdown, comp off since 3/10/2001.
|
| ANS
| anolla@naic.edu
| 4/4/2001
| Dewar turned off to replace crosshead. Also Dewar will be taken down to check
for leaks.
|
| ANS
| anolla@naic.edu
| 4/20/2001
| Dewar was tested for leaks with the Varian leak tester, no leak was found. The
crosshead needs to be cleaned out.
|
| ANS
| anolla@naic.edu
| 5/23/2001
| A/L sw at the radar room bypassed , always on antenna except when 430 mhz
xmmter up at CH 1.
|
| ANS
| anolla@naic.edu
| 7/27/2001
| The recvr was installed back, but there is not ant. System with 50 ohms at input.
|
| ANS
| anolla@naic.edu
| 8/15/2001
| The 327 kHz antenna was installed; it has new hybrids on it ( TRM ) .Also the
match xformers were removed from it.
|
| ANS
| anolla@naic.edu
| 8/21/2001
| The channel A A/L sw is bad . Also this morning due to a power glitch the cryo
compressor tripped off.
|
| ANS
| anolla@naic.edu
| 8/22/2001
| System is cold. Just VP to tell, problems with the Lakeshore reading. JR is
working with the A/L driver.
|
| ANS
| anolla@naic.edu
| 8/23/2001
| The A/L driver was ok , the problem was the 24 vdc power supply which drives
it had a 7 v ripple that was causing the A/L relays to chattering.
|
| ANS
| anolla@naic.edu
| 8/27/2001
| The cryo comp replaced , it was ED 007 (back to WSB) now is ED 016. Also this
morning the comp was found tripped off , maybe due to power dips.
|
| ANS
| anolla@naic.edu
| 8/30/2001
| The system has been cooling down because it warmed up this morning when we
try to switch cryo compressors that didn't worked out , comp ED 011 still has problems.
|
| ANS
| anolla@naic.edu
| 9/20/2001
| System been cooled with compressor ED 021 ( Austin 125 ).It had a leak on it,
but it is not in its cryo lines , we check them and they are ok.
|
| ANS
| anolla@naic.edu
| 9/20/2001
| System cooled. Need Tsys check out. Also the problem with its cryo reading fixed
by itself , what it was ? we don't know.
|
| ANS
| anolla@naic.edu
| 9/27/2001
| System again hot , problems with the cryo. System at room temp.
|
| ANS
| anolla@naic.edu
| 10/18/2001
| Jose Rosa modified the Lakeshore ckt in order to be able to take downstairs
readings . Also WG found a problem with the cryo reading board ( loose Vcc cable on it).
Both chassis were installed back at the Dome.
|
| ANS
| anolla@naic.edu
| 10/21/2001
| 327 mhz protected from the SB xmitter thru the CB/SBW shutter control relay (
if shutter is closed it will send the 327 mhz to load, so have to be aware of that )
|
| ANS
| anolla@naic.edu
| 1/30/2002
| Chan A amp is sn 350-12 and its biases are:
First stage:Vd 1.42 vdc,Id 3.7 ma
Second stage:Vd 1.44 vdc,Id 4.9 ma. Chan B amp is sn 350-14 and its biases are:
First stage:Vd 1.51 vdc, Id 4.3 ma
Second stage: Vd 1.51 vdc, Id 5.0 ma.
|
| ANS
| anolla@naic.edu
| 4/1/2002
| Spare amp 350-13 came fixed from NRAO.
|
| ANS
| anolla@naic.edu
| 7/31/2002
| New receiver at room temp installed.
|
| LW
| lwray@naic.edu
| 8/5/02
| 2 MHz comb that was seen in the band has been traced to the radar box. The box has
been re-shielded and re-installed in the dome. Data was taken today with the box ON.
The 2 MHz comb is still there, but VERY much reduced in power.
|
| ans | anolla@naic.edu | Tuesday, October 29, 2002 11:32:29 | New Angle Linear amps on it , model 2X327GST (room temp).Channel A is S/N 11467 & Channel B is S/N 11468 .Spare is S/N 11469.They are 20º better than the former ones.
|
| LSWl | lwray@naic.edu | Monday, November 4, 2002 15:01:33 | Tested LNA box using NFA, found 65K with 5 dB ENR, and ~55K with 15 dB ENR. Tsys program reads 110K at 9am. (both channels read ~same in these cases) At 2pm, Tsys program reads 150K. RFI environment is the cause of the increase.
|
| ansa | anolla@naic edu. | Friday, December 6, 2002 13:02:39 | Connectors added to the antenna cables to protect them from been broken.
|
| lsw | lwray@naic.edu | Thursday, January 22, 2004 16:08:35 | Jan 7: receiver and feed moved.
|
| lsw | lwray@naic.edu | Thursday, January 22, 2004 16:09:44 | Jan 21: feed hooked up (thanks to PP) and receiver functioning in its new spot on the rotary floor (next to X Band)
|
| lswl | lwray@naic.edu | Thursday, July 1, 2004 15:08:19 | Calixto and Carlos testing for RFI today.
|
| ans | anolla@naic.edu | Thursday, September 30, 2004 13:34:56 | When installing the hybrid for the polarization test of the system we found sort a few things , first : channel A antenna connector was unsolder , several sma connectors loose at the amp chassis. The Tsys with system polarized were ; chann A =119 deg , chann B=120 deg.
|
| ans | anolla@naic.edu | Monday, October 4, 2004 13:06:31 | Tsys was +/- 20 K deg higher at chann B than chann A. Leading cables from the antenna to the recvr were unsolder. After fixing the cables both channels were at 120 K deg
|
| ans | anolla@naic.edu | Monday, October 11, 2004 12:38:16 | System been polarized with Quad Hybrid.
|
| ans | anolla@naic.edu | Tuesday, October 12, 2004 13:02:48 | System been kept polarized with quad hybrid.
|
| test | test | Wednesday, March 30, 2005 14:41:02 | test
|
| W. Genter | wgenter@naic.edu | Wednesday, September 7, 2005 09:42:28 | There has been a 16KHz comb in the bandpass of this receiver since May 2005. This did not affect pulsar work, but for all others this comb prevented observations in this band. Several efforts were made to find the source of this without success. In late August 2005, Rey Velez, PP and Myself found the source of the interference. The source is the Alfa rotation control motor. The Motor control cable from to the rotation control motor was not grounded to the chassis. At this time, the Alfa rotation drive motor must have power removed to observe at 327MHz.
|
| ans | anolla@naic.edu | Tuesday, November 29, 2005 10:53:04 | The 300 Mhz HP filters for RFI didn't work as expected.No change.
|
| ans | anolla@ | Wednesday, November 30, 2005 13:27:20 | The 300 Mhz HP filters were removed.Didn't work properly.
|
| ans | anolla@naic.edu | Thursday, April 13, 2006 13:02:21 | The rf cable of channel A between the front end and the filter at the post amp was replaced , was bad. Also the cal cable of channel B between post amps and the antenna was replaced.
|
| ans | anolla@naic.edu | Monday, May 15, 2006 14:45:30 | The missing NRAO cooled amps were found in a box at the cryo shop.Back to 11F-K04-1 drawer in the stock room.
|
| W. Genter | wgenter@naic.edu | Monday, July 24, 2006 10:55:35 | The new Dewar is in process of being finished this date. The heat shields are being plated. Dewar (no paint) directional couplers have been received and the cryo amplifiers located.
|
| W. Genter | wgenter@naic.edu | Wednesday, August 16, 2006 07:43:18 | The new dewar has arrived.
|
| W. Genter | wgenter@naic.edu | Thursday, September 14, 2006 09:42:24 | Two DCI cavity filters have been received for this receiver. The filters are centered at 327MHz with a BW of 50MHz.
|
| ans | anolla@naic.edu | Friday, November 17, 2006 11:35:17 | Post amp chassis down for the filter bank installation.System down until further notice.
|
| ans | anolla@naic.edu | Wednesday, November 22, 2006 16:11:50 | 327 Mhz back to service.New filter bank on it , but can't swap filters yet.Set to 327Mhz / 30 Mhz B.W.
|
| ans | anolla@naic.edu | Wednesday, November 29, 2006 14:23:18 | Filter bank control installed.Is bit 402 of the HP3488 I. The 50 Mhz B.W. filter is the default one.
|
| ans | anolla@naic.edu | Monday, December 11, 2006 12:51:07 | The receiver configuration was reset in order to have the shortest rf cables between the DCI cavity filters & the post amp.So the DCI filters are at the bottom of the rf cabinet with the post amp , the recvr control was removed to the middle of the rf cab (open frame).
|
| ans | anolla@naic.edu | Friday, December 22, 2006 08:59:14 | Due to the new configuration with the filter bank , the system gain was increased by 6 db when using the 50 Mhz B.W. filters .So a fix 6 db pad was added to the output of the rf of it to reduce these gain.Gain came from 3 db from doubling the B.W. and 3 more db from less loss of the 50 Mhz B.W. filter.
|
| ans | anolla@naic.edu | Monday, April 23, 2007 09:53:04 | System down due to painting project.
|
| crv | | Friday, June 6, 2008 10:45:25 | On friday May the 30th the new dewar was installed and the chasis were relocated.
|
| crv | | Friday, June 6, 2008 10:47:49 | Tuesday June the 3rd the dewar amp were tested an the cryo temp checked finding problems with Cha Stage 2 amp ccurrent.
|
| crv | | Friday, June 6, 2008 10:50:05 | Wednesday june the 4th the dewar was open and the problem was repaired. power connector over thighten. The connector was lossen and thight corectly.
|
| crv | | Tuesday, June 10, 2008 10:24:17 | Today the new dewar as started cooldown with a small compresssor ed 021.
|
| crv | | Wednesday, June 18, 2008 16:00:40 | On friday june the 12th the reciever being used is the coolreciever.
|
| crv | | Wednesday, June 18, 2008 16:04:42 | 50 MHz filter have been mounted over the feed horn and attached under the floor in brackets.
|
| CRV | | Monday, October 13, 2008 14:16:11 | On friday october the 10th the crosshead stop rotating.
Motor was stucked in one position. Crosshead ED 20 was removed Large amount of water found on both stage displacers. Dewar cylinder was cleaned and dried and a brand new frig was installed.
|
| Stephen Jensen | sjensen@naic.edu | Thursday, December 11, 2008 16:32:49 | 4:00 PM: The 327 receiver is warming up nicely.
|
| crv | | Thursday, August 6, 2009 15:10:26 | reciever was vacuumed and is cooled down today 8/6/2009 after sunday night power failure and wednesday power failure .
|
| CRV | | Monday, January 4, 2010 16:46:51 | System was removed from Comp. ED015 and connected to ALFA'S new compressor C2 . This compressor has now three cryo pumps (alfa 2, 327MHz , XB).
|
| crv | | Wednesday, January 13, 2010 10:09:32 | System was warming up on 1/12/10 70k = 165k. The system was totally warmed up vac.pump connected and will be coolled on 1/13/10.
|
| crv | | Monday, February 22, 2010 14:16:38 | system is now connected to ALFA C1 compressor.
|
| crios | | Friday, April 30, 2010 17:24:02 | system was disconnected from Compressor C2 and conected on the spare comp. ED015 to allow conecting Alfa second cryo pump to the C1 compressor. This change was made due to C2 comp. failing.
|
| crios | | Friday, April 30, 2010 17:24:58 | The change of compressor was on April 23rd
|
| crv | | Wednesday, June 2, 2010 10:41:14 | The system is down due to an internal broken line on the compressor C3. Below the main dish is still flooded and is dificult to lower or raise a compressor.
|
| crv | | Monday, September 27, 2010 17:02:59 | system was warmed and purged an recooled is operational.
|
| GR | grivera@naic.edu | Tuesday, November 30, 2010 16:22:37 | The dewar was warmed up yesterday (Nov.29). We started the cooldown today.
|
| crv | crios@naic.edu | Monday, December 6, 2010 13:57:33 | Reciever is been warmed to purge and do vac. to the dewar. This is do to the power failure of last week.The reciever will be ready for dec-9-2010 observation.
|
| crv | crios | Wednesday, December 8, 2010 13:47:04 | The cryo pump was cleaned and purged. The dewar vac. was o.k. and the system started cooldown.
|
| crv | crios@naic.edu | Thursday, December 9, 2010 10:06:49 | System is cold and the vac valve was closed.
|
| crv | | Tuesday, March 1, 2011 11:48:06 | The cryopump was replaced on friday 2/28/2011.
ED 018 was installed system is cold and vac.pump was removed today.
|
| crios | | Monday, May 23, 2011 16:16:19 | The system was sounding contaminated the dewar was partially warmed up and purged. The system is already cold.
|
| crios | | Monday, June 6, 2011 14:00:56 | On friday june the 3rd a hybrid was installed on the feed horn . The Rf cables from the the feed were broken on the base and were re-soldered.
|
| Carlos Rios | crios@naic.edu | Thursday, February 2, 2012 10:15:49 | The cooling system had a leak and a flexible line on the transmitter room was replaced. After replacing the line the XB and 327 Mhz were contaminated. The system was turned off on Feb. 1st 2012. Started cooldown today.
|
| Carlos Rios | crios@naic.edu | Thursday, February 2, 2012 10:25:49 | During mid Dic. 2011 a hybrid was connected to the 327 feed antena. the connection is still in place.
|
| crios | crios@naic.edu | Monday, May 7, 2012 10:14:17 | The cryo pump sounded contaminated. The system was turned off.
|
| Denis Urbain | durbain@naic.edu | Wednesday, April 3, 2013 13:15:57 | Compressor lost pressure during the weekend. The receiver was warming up because of the lost cooling capacity. The cold head was stopped to restore the vacuum. After 12 hours the vacuum was good and the cold head started again. Until now the compressor has kept is pressure, no clear explanation for the loss of pressure during the weekend.
|